Use the fundamental theorem ofhomomorphism for groups to prove


1
Expert's answer
2021-05-12T02:59:46-0400

 zassehaus (buttergly lemma)


statement:- let H,K be subgroups of a group G. let H',K' be normal subgroup of H and K, respectively, then


(i) (H∩K')H' ⊲ (H∩K)H'


(ii) (H'∩K)K' ⊲ (H∩K)K'


(iii)(H∩K)H'/ (H∩K')H' ≅ (H∩K)K'/(H'∩K)K'


proof:- since H,K are subgroup of G, (H∩K) is also a subgroup of G. Moreover, (H∩K)⊂K, so (H∩K) is a subgroup of K also.


No K'is normal (H∩K) = (H∩K)∩K'is a normal subgroup d (H∩K). Similarly, (H'∩K) is normal in (H∩K). 

Consequently product of two normal subgroups,(H∩K') (H'∩K)= L, is a normal subgroup of (H∩K). 

Now, (H∩K)CH,H'<H implies that (H∩K) H' is a subgroup of H. Similarly, (H∩K)K' is a subgroup of K.



Define a map


ϕ: (H∩K)H'→ (H∩K)/L


by


ϕ(xy) = xL,


x∈H∩K,y∈H'


ϕ is well defined, since for x,x1 ∈H∩K,y,y1∈H'




if


xy =x1y1


x1-1x=y1y-1∈(H∩K)∩H'= (H'∩K) 


x1-1x∈(H'∩K)⊂L


x1-1x∈L


xL=x1L


ϕ (xy) =ϕ (x1y1)




 ϕ is clearly onto.


Now for x,x1 ∈H∩K,y,y'∈H'


 ϕ [(xy)(x1y1)]= ϕ [xx1(x1-1yx1)y1]= ϕ [xx1y2y1]


                          where y2 =x1-1yx1∈H'


=xx1

=xLx1

=ϕ (xy) ϕ (x1y1)


so ϕ is an epimorphism. 


Further, 




xy ∈ker ϕ ⟺ϕ (xy) = L 

         ⟺xL =L 

         ⟺x∈L 

         ⟺x=x1y1, with x1 ∈H∩K', y1 ∈H'∩K


thus xy ∈ker ϕ if only if 

xy=(x1y1)y

   =x1(y1y) ∈(H∩K')H'


Therefore, ker ϕ = (H∩K') H'. 




This implies that (H∩K')H'⊲ (H∩K)H'. 




Finally, by Fundamental Homomorphism Theorem, 




we have (H∩K)H'/(H ∩K')H'≅ (H∩K)/L. 




A symmetric argument shows that 




(H'∩K)K'⊲ (H∩K)K', and (H∩K)K'/(H'∩K)K'≅(H∩ K)/L. 




Consequently, (H∩ K) H'/(H∩K')H'≅ (H∩K) K'/(H'∩K) K'.